Leonardo, who is married but files separately, earns $150,000 of taxable income. He also has $13,250 in city of Tulsa bonds. His wife, Theresa, earns $74,000 of taxable income.

If Leonardo and his wife file married filing jointly in 2019, What is Leonardo and Theresa's effective tax rate for 2019 (rounded)? (Use tax rate schedule)

24.67 percent

19.37 percent

25.74 percent

15.74 percent

17.75 percent
Answer #1
Total taxable income 224000 =150000+74000
Tax on $224000 42109 =28765+(224000-168400)*24%
Tax payable 42109
Divide by Total Income 237250 =224000+13250
Effective tax rate for 2019 17.75%
Option E 17.75 percent is correct